(g) Prepare Journal Entries for Notes Receivable Akron Corporation engaged in the following transactions involving promissoryJournalize these transactions in the journal provided. Explanations are not needed. Akron records adjusting entries monthly.
Answer #1

1. General Journal (Amounts in $)

Date Accounts Title and Description Debit Credit
20X1
Sept 1 Notes Receivable 160,000
Land 160,000
(To record sale of land)
Nov 1 Notes Receivable 4,000
Accounts Receivable 4,000
(To record notes received from Fred Hansen)
Dec 1 Accounts Receivable 4,000
Notes Receivable 4,000
(To record dishonored Note)
Dec 31 Interest Receivable a 933
Interest Revenue 933
(To record accrued Interest)
20X2
March 1 Cash 165,600
Notes Receivable 160,000
Interest Revenue b 5,600
(To record the amount received from note)

Working Notes:

a) Calculation of Interest for the month of December =(160,000*7%)/12

= $11,200/12 = $933

b) Calculation of Interest Revenue from September to December = (160,000*7%)/2

=$5,600